OpenStudy (anonymous):

Polar coordinates. Find the area of the region that lies inside the first curve and outside the second curve. first curve: r^2=8cos(2x) second surve: r=2 i mainly need help finding the limits of integration to this problem. i set 4=8cos(2x) cos2x=1/2 x = pi/3 and -pi/3(first and fourth quadrant is where cos is positive.) but since there is a 2 in the cos, i divide both by 2 and get pi/6 as upper limit and -pi/6 as bottom limit of the integral. but when i integrate, i still get the wrong answer. I have even used an integral calculator to see if i was integrating properly. (i was)

OpenStudy (anonymous):

\[.5\int\limits_{-\pi/6}^{\pi/6}8\cos(2x)-(2)^2\]

OpenStudy (anonymous):

correct answer is 2.74

OpenStudy (anonymous):

Thanks!!